# Temperature dependent ultracold neutron transmission in D2 gas -- a test of the Young-Koppel model

### v1

@@ I Introduction
The present work investigates the interaction between slow neutrons and individual deuterium molecules. This interaction, by its fundamental nature, has intrigued researchers from the early days of neutron physics [8, 9, 10] and remains highly topical until today in nuclear physics, for example, in the design of neutron sources [11]. It also has important practical implications for the non-standard operating conditions of the PSI UCN source when considerable D2 vapor pressure is present above the solid D2 [12], as UCN-D2 scattering can lead to significant UCN losses via up-scattering - i.e. neutrons acquire large enough energies to leave the UCN energy range.
In this work we tested the Young-Koppel model [13] that can be applied to describe the interaction between UCNs and H2 or D2 gases. It is the first time that the Young-Koppel model was tested over a wide temperature range from about 100 K to above 300 K.
@@ II Theoretical description
A significant time later, the total cross section of deuterium gas at 25 K was measured [19]. However, the YK model and a low energy approximation model could not be distinguished at this single low temperature, thus it did not provide a full corroboration of the YK model. New gas measurements were performed at room temperature [20, 21], where the authors scaled with $T^{1/2}$ the 25 K-value reported in Ref. [19] to an RT value. With simple reasoning, in the coordinate system of the molecules, the UCNs propagate with the mean velocity of the Maxwell-Boltzmann gas, $w\propto T^{1/2}$, and the interaction rate $\sigma w$ is invariant to changing the coordinate system (see also Eq. (4) in Ref. [19]). This assumption of the low-energy low-temperature approximation (LETA) is over-simplified at higher D2 gas temperatures. The LETA model is similar to the monoatomic gas model [22], however, it uses an effective scattering cross section neglecting rotations and spin-correlation effects. The YK theory shows an important deviation of the up-scattering cross section from a pure $\sqrt{T}$ dependence, which will be demonstrated experimentally in the present paper. Molecular vibrations are also included in the YK model, however, they are negligible in our experimental temperature range (vibration modes are populated only above 4000 K).
@@ III.1 Experimental setup
The gas cell was sealed on both sides with 0.1 mm thick aluminum (AlMg3) windows to separate the gas volume from the vacuum in the UCN guides. This window was supported by stainless-steel grids that covered approximately 10 % of the tubes cross section. The gas cell was made of stainless steel as a compromise between temperature stability and UCN guiding properties. Cooling was done with a cold gas system where the cold N2 is guided first around the center of the cell and then in parallel to both ends. To achieve a good temperature equilibration with the windows, it was important that the cooling gas also cooled the stainless-steel UCN guides connected to both sides of the cell. Calculations and finite element simulations showed that the highest temperature gradient was expected starting from the center of the window (“hot”) to the point where the cold N2 gas is first reaching the cell. Therefore, we positioned several temperature sensors inside the gas volume which were introduced from the top via a pipe at a 45-degree angle to the horizontal. Additional temperature sensors were used on the outside of the cell on both aluminum windows. The $T$-sensors were calibrated beforehand with liquid N2 and distilled ice water. In the final analysis, a time- and position-averaged temperature value of the sensor readings was used. We also checked, that the impact of temperature gradients on the final result is negligible compared to our main systematic uncertainty in the UCN dwell-time spectrum. Pressure sensors and the gas inlet were connected by a separate flange. We also made sure that there was no relevant effect of a temperature gradient on the pressure measurement.
@@ III.2 Analysis concept
$\tau=(\rho\cdot\sigma\cdot v_{\mathrm{n}})^{-1}=\left(\frac{p}{k_{\text{B}}\sqrt{T}}\epsilon\right)^{-1},$ (3)
@@
$\epsilon=\frac{\sigma\cdot v_{\mathrm{n}}}{\sqrt{T}}$ (4)
@@
$P_{\mathrm{gas}}^{v_{z}}(l,p,T;\frac{\epsilon}{v_{z}})=\exp{\left(-\frac{lp}{k_{\text{B}}\sqrt{T}}\cdot\frac{\epsilon}{v_{z}}\right)},$ (5)
@@
$\epsilon(T)=\epsilon_{0}\cdot(\epsilon_{a}-\epsilon_{T}\cdot T),$ (6)
where $\epsilon_{0}$ represents the LETA approximation, and $\epsilon_{a}=0.98\pm 0.03$ and $\epsilon_{T}=(6.26\pm 0.37)\cdot 10^{-4}$ K-1. Besides the fit error, there is a $\sim 1~{}\%$ error from the numerical uncertainty of the MC integration and $\sim 2~{}\%$ from uncertainty in the input parameters like the equilibrium distance of the nuclei. In total, we estimate the slope error to be below $6~{}\%$.
Other than in the theory, in the analysis of the measurements, one cannot disentangle $\epsilon_{a}$ from $\epsilon_{0}$. However, we were looking for the temperature dependence additional to $\sqrt{T}$. Thus, the signature we were searching for was covered by $\epsilon_{T}$ and we thus set $\epsilon_{a}$ to the theoretically expected value of $0.98$ valid in our temperature range of interest.
@@
$P_{\mathrm{gas}}(l,p,T;\epsilon)=\int P_{\mathrm{gas}}^{v_{z}}(l,p,T;\frac{\epsilon}{v_{z}})\cdot f(v_{z})\mathrm{d}{v_{z}},$ (7)
@@
We extensively investigated the measured data sets of two gases: neon (Ne) to serve as a test candidate for the LETA model and D2 for the YK model. We checked whether $\epsilon_{T}$ is compatible with 00 in the case of Ne, and whether for D2 it is compatible with the value $(6.26\pm 0.37)\cdot 10^{-4}\,\mathrm{K}^{-1}$ from the theoretical Young-Koppel prediction.
@@ IV Systematics
$P_{\mathrm{gas}}(T)=\frac{1}{R\cdot P_{\mathrm{window}}(T)}\cdot\frac{N_{A}(p,T)}{N_{B}},$ (8)
@@ IV.1 Reference value of the setup with empty cell
$R=\frac{N_{A,\mathrm{ref}}(p=0,T_{\mathrm{ref}})}{P_{\mathrm{window}}(T_{\mathrm{ref}})}\cdot\frac{1}{N_{B,\mathrm{ref}}}$ (9)
@@ IV.2 Temperature dependence for the empty cell
$P_{\mathrm{window}}(T)=\frac{C_{T}\cdot T+C_{0}}{C_{0}},$ (10)
where $C_{T}=(-1.73\pm 0.03)\cdot 10^{-3}K^{-1}$ and $C_{0}=5.470\pm 0.006$. Since it is impossible to disentangle the absolute transmission probability of the setup, the function was built in a way that it is $1$ at $0\,\mathrm{K}$. This is valid since the absolute scale will cancel out in $P_{\mathrm{gas}}$ with the absolute scale of $R(T_{\mathrm{ref}})$. The relevant part for this investigation is, that the empty cell transmission changes with the temperature so that the transmission probability of the empty cell at around $100\,\mathrm{K}$ is about $6\%$ higher than at RT. This is plausible considering up-scattering on phonons in the aluminum window causing a previously observed temperature dependence of the total UCN cross section on aluminum [27, 28].
@@ IV.3 Pressure-induced change of cell length
$l=127.6\,\mathrm{mm}+p\cdot 1.060\cdot\mathrm{mm/bar}$ (11)
@@ IV.4 Calibration of the dwell-time spectrum
$N_{\mathrm{stored}}=a_{storage}\cdot\exp{(-t_{\mathrm{dwell}}/\tau_{storage})}$ (14)
@@ IV.5 Systematic test with reference gases
Again, accounting for the correct dwell-time spectrum including stored UCNs we found good agreement with the literature and calculated the values shown in Tab. 1. For reference, we considered 6.6 m/s UCN-velocity in order to compare to previous work [20] applying the same textbook equation (Eq. (11) in [20] from Ref. [29]) to convert from free atoms at rest to a Maxwell-Boltzmann gas. The values for the total cross section were obtained by fitting Eq. 7 to transmission data using the dwell-time spectrum from Fig. 3. The statistical uncertainty is extracted from the fit. The systematic uncertainty was obtained by maximum error propagation by fitting with minimum and maximum values derived from the uncertainties on $\tau_{\mathrm{storage}}$ and $a_{\mathrm{storage}}$ (Eq. 15).
@@ V Results and discussion
The main transmission measurements were performed with gaseous D2 and Ne in the pressure range from 0 to 1 bar, and from $110\,\mathrm{K}$ to $365\,\mathrm{K}$. The pressure, $p$, and the temperature, $T$, of the gas are measured directly with sensors in each measurement.
In Eq. 5 the exponent of the transmission probability of a neutron is given as ${\left(-\frac{lp}{k_{\text{B}}\sqrt{T}}\cdot\frac{\epsilon}{v_{z}}\right)}$, where the cell length, $l$, is dependent on the absolute pressure, $p$, according to Eq. 11. Consequently, the transmission can be plotted as a function of ${\left(\frac{lp}{\sqrt{T}}\right)}$. By this choice, one expects a close-to-exponential dependence of the measured transmissions on ${\left(\frac{lp}{\sqrt{T}}\right)}$, which also includes the broadening due to the dwell time spectrum. These transmission profiles are seen in the top plots of Fig. 4 for D2 gas and Fig. 5 for Ne.
@@
Figure 4: Top: UCN transmission data (black dots) in D2 gas as a function of ${\left(\frac{lp}{\sqrt{T}}\right)}$ (at all measured $T$-s), and the best fits of the two alternative models with example values for the independent variable $T$ in case of YK (blue=100 K, yellow=300 K, red=400 K lines), and one line with LETA (black). Center: Point-by-point residuals of the YK and LETA model fit to the measured data, showing data at all measured temperatures in a color scale. Bottom: Histograms of the above residuals showing a much narrower and symmetrical distribution for the case of the YK model.
@@
In contrast to D2, the Ne data fits with different temperatures shown in Fig. 5 converge with a slope in $T$ compatible with zero, see also Tab. 2, hence the fit involving Eq. 6 reproduces the LETA model.
@@
The fit results are summarized in Tab. 2. As in the results for other gases (Tab. 1), the uncertainties of $\epsilon_{0}$ are for both gases, D2 and Ne, dominated by the systematic contribution from $N_{\mathrm{dwell}}$, whereas, for $\epsilon_{T}$ the uncertainties are dominated by statistics, as the latter is a relative parameter. The neutron counts per measurement point for both detectors were above $3\cdot 10^{5}$. However, considering only pure neutron count rate statistics would neglect other effects, such as minor changes in the neutron spectrum due to slow changes in the neutron production efficiency of the source.
@@
species $\epsilon_{0}$ stat. syst. $\epsilon_{T}$ stat. syst. effective $\epsilon$
(barn $(\frac{\text{barn }m/s}{K^{1/2}})$$\pm$ $\pm$ $10^{-6}/K$$\pm$ $\pm$ (barn $(\frac{\text{barn }m/s}{K^{1/2}})$
Ne 100.6 1.0 10.0 54 38 10 97.0
D2 1238.8 3.8 54.8 559 10 6 1006.3
@@
Returning to Tab. 1, the absolute value of the D2 total cross section for UCNs is given under ‘literature’ as calculated by integrating Eqs. (A4) and (A5) in Ref. [13] and adding in a 2:1 ortho-para RT-equilibrium ratio. We also measured the total cross section of UCNs in H2 gas. The data were analyzed by applying the Young-Koppel model, however, in a more limited temperature range (250-350 K) and thus with less sensitivity to the details of this model. The theoretical YK differential cross sections, Eqs. (29)-(30) in Ref. [13], for para- and ortho-hydrogen were integrated over final energy and solid angle, and added in a 1:3 para to ortho equilibrium ratio at RT. The measurement result for H2 and the theoretical expectation are in very good agreement, see Tab. 1, consistent within the systematic uncertainties.

### v4

@@ I Introduction
The present work investigates the interaction between slow neutrons and individual deuterium molecules. This interaction, by its fundamental nature, has intrigued researchers from the early days of neutron physics [8, 9, 10] and remains highly topical until today in nuclear physics, for example, in high precision measurements of the hadronic interaction [11, 12] and in the design of neutron sources [13, 14]. It also has important practical implications for the non-standard operating conditions of the PSI UCN source when considerable D2 vapor pressure is present above the solid D2 [15], as UCN-D2 scattering can lead to significant UCN losses via up-scattering - i.e. neutrons acquire large enough energies to leave the UCN energy range.
In this work we tested the Young-Koppel model [16] that can be applied to describe the interaction between UCNs and H2 or D2 gases. It is the first time that the Young-Koppel model was tested over a wide temperature range from about 100 K to above 300 K. The temperature dependence is its unique signature for distinguishing it from the monoatomic approximation.
@@ II Theoretical description
A significant time later, the total cross section of deuterium gas at 25 K was measured [22]. However, the YK model and a low energy approximation model could not be distinguished at this single low temperature, thus it did not provide a full corroboration of the YK model. New gas measurements were performed at room temperature [23, 24], where the authors scaled with $\sqrt{T}$ the 25 K-value reported in Ref. [22] to an RT value. With simple reasoning, in the coordinate system of the molecules, the UCNs propagate with the mean velocity of the Maxwell-Boltzmann gas, $w\propto\sqrt{T}$, and the interaction rate $\sigma w$ is invariant to changing the coordinate system (see also Eq. (4) in Ref. [22]). This assumption of the low-energy low-temperature approximation (LETA) is over-simplified at higher D2 gas temperatures. The LETA model is similar to the monoatomic gas model [25], however, it uses an effective scattering cross section neglecting rotations and spin-correlation effects. The YK theory shows an important deviation of the up-scattering cross section from a pure $\sqrt{T}$ dependence, which will be demonstrated experimentally in the present paper. Molecular vibrations are also included in the YK model, however, they are negligible in our experimental temperature range (vibration modes are populated only above 4000 K).
@@ III.1 Experimental setup
The gas cell was sealed on both sides with 0.1 mm thick aluminum (AlMg3) windows to separate the gas volume from the vacuum in the UCN guides. This window was supported by stainless-steel grids that covered approximately 10 % of the tubes cross section. The gas cell was made of stainless steel as a compromise between temperature stability and UCN guiding properties. Cooling was done with a cold gas system where the cold N2 is guided first around the center of the cell and then in parallel to both ends. To achieve a good temperature equilibration with the windows, it was important that the cooling gas also cooled the stainless-steel UCN guides connected to both sides of the cell. Calculations and finite element simulations showed that the highest temperature gradient was expected starting from the center of the window to the point where the cold N2 gas is first reaching the cell. Therefore, we positioned several temperature sensors inside the gas volume which were introduced from the top via a pipe at a 45-degree angle to the horizontal. Additional temperature sensors were used on the outside of the cell on both aluminum windows. The $T$-sensors were calibrated beforehand with liquid N2 and distilled ice water. In the final analysis, a time- and position-averaged temperature value of the sensor readings was used. We also checked, that the impact of temperature gradients on the final result is negligible compared to our main systematic uncertainty in the UCN dwell-time spectrum. Pressure sensors and the gas inlet were connected by a separate flange. For the pressure sensors a two point calibration was done at atmospheric pressure (using meteorologic reference data) and after evacuation with a two stage pump train (both with a systematic error smaller than 0.1%). We also made sure that there was no relevant effect of a temperature gradient on the pressure measurement.
@@ III.2 Analysis concept
$\tau=(\rho\text{ }\sigma\text{ }v_{\mathrm{n}})^{-1}=\left(\frac{p}{k_{\text{B}}\sqrt{T}}\text{ }\epsilon\right)^{-1},$ (3)
@@
$\epsilon=\frac{\sigma\text{ }v_{\mathrm{n}}}{\sqrt{T}}$ (4)
@@
$P_{\mathrm{gas}}^{v_{z}}(l,p,T;\frac{\epsilon}{v_{z}})=\exp{\left(-\frac{lp}{k_{\text{B}}\sqrt{T}}\text{ }\frac{\epsilon}{v_{z}}\right)},$ (5)
@@
$\epsilon(T)=\epsilon_{0}\text{ }(\epsilon_{a}-\epsilon_{T}\text{ }T),$ (6)
where $\epsilon_{0}$ represents the LETA approximation, and $\epsilon_{a}=0.98\pm 0.03$ and $\epsilon_{T}=(6.26\pm 0.37)\times 10^{-4}$ K-1. Besides the fit error, there is a $\sim 1~{}\%$ error from the numerical uncertainty of the MC integration and $\sim 2~{}\%$ from uncertainty in the input parameters like the equilibrium distance of the nuclei. In total, we estimate the slope error to be below $6~{}\%$.
Other than in the theory, in the analysis of the measurements, one cannot disentangle $\epsilon_{a}$ from $\epsilon_{0}$. However, we were looking for the temperature dependence additional to $\sqrt{T}$. The signature we were searching for was covered by $\epsilon_{T}$. We thus set $\epsilon_{a}$ to the theoretically expected value of $0.98$ valid in our temperature range of interest, and propagate its uncertainty into the final systematic error.
@@
$P_{\mathrm{gas}}(l,p,T;\epsilon)=\int P_{\mathrm{gas}}^{v_{z}}(l,p,T;\frac{\epsilon}{v_{z}})\text{ }f(v_{z})\mathrm{d}{v_{z}},$ (7)
@@
We extensively investigated the measured data sets of two gases: neon (Ne) to serve as a test candidate for the LETA model and D2 for the YK model. We checked whether $\epsilon_{T}$ is compatible with 00 in the case of Ne, and whether for D2 it is compatible with the value $(6.26\pm 0.37)\times 10^{-4}\,\mathrm{K}^{-1}$ from the theoretical Young-Koppel prediction.
@@ IV Systematics
$P_{\mathrm{gas}}(T)=\frac{1}{R\text{ }P_{\mathrm{window}}(T)}\text{ }\frac{N_{A}(p,T)}{N_{B}},$ (8)
@@ IV.1 Reference value of the setup with empty cell
$R=\frac{N_{A,\mathrm{ref}}(p=0,T_{\mathrm{ref}})}{P_{\mathrm{window}}(T_{\mathrm{ref}})}\text{ }\frac{1}{N_{B,\mathrm{ref}}}$ (9)
@@ IV.2 Temperature dependence for the empty cell
$P_{\mathrm{window}}(T)=\frac{C_{T}\text{ }T+C_{0}}{C_{0}},$ (10)
where $C_{T}=(-1.73\pm 0.03)\times 10^{-3}K^{-1}$ and $C_{0}=5.470\pm 0.006$. Since it is impossible to disentangle the absolute transmission probability of the setup, the function was built in a way that it is $1$ at $0\,\mathrm{K}$. This is valid since the absolute scale will cancel out in $P_{\mathrm{gas}}$ with the absolute scale of $R(T_{\mathrm{ref}})$. The relevant part for this investigation is, that the empty cell transmission changes with the temperature so that the transmission probability of the empty cell at around $100\,\mathrm{K}$ is about $6~{}\%$ higher than at RT. This is plausible considering up-scattering on phonons in the aluminum window causing a previously observed temperature dependence of the total UCN cross section on aluminum [30, 31].
@@ IV.3 Pressure-induced change of cell length
$l=127.6\,\mathrm{mm}+p\times 1.060~{}\mathrm{mm/bar}$ (11)
@@ IV.4 Calibration of the dwell-time spectrum
$N_{\mathrm{stored}}=a_{storage}\text{ }\exp{(-t_{\mathrm{dwell}}/\tau_{storage})}$ (14)
@@ IV.5 Systematic test with reference gases
Again, accounting for the correct dwell-time spectrum including stored UCNs we found good agreement with the literature and calculated the values shown in Table 1. We also obtained very good agreement with the measurements of Ne, Ar, Xe, CF4 in [23] at RT, the difference being smaller than the error bars. For reference, we considered 6.6 m/s UCN-velocity in order to compare to previous work [23] applying the same textbook equation (Eq. (11) in [23] from Ref. [32]) to convert from free atoms at rest to a Maxwell-Boltzmann gas. The values for the total cross section were obtained by fitting Eq. 7 to transmission data using the dwell-time spectrum from Fig. 3. The statistical uncertainty is extracted from the fit. The systematic uncertainty was calculated using maximum error propagation by fitting with minimum and maximum values derived from the uncertainties on $\tau_{\mathrm{storage}}$ and $a_{\mathrm{storage}}$ (Eq. 15).
@@ V Results and discussion
The main transmission measurements were performed with gaseous D2 (99.6 % D2 purity with a fraction of HD molecules of 0.4 % which has a negligible contribution compared to the main systematic effects discussed) and Ne in the pressure range from 0 to 1 bar, and from $110\,\mathrm{K}$ to $365\,\mathrm{K}$. The pressure, $p$, and the temperature, $T$, of the gas are measured directly with sensors in each measurement.
In Eq. 5 the exponent of the transmission probability of a neutron is given as ${\left(-\frac{lp}{k_{\text{B}}\sqrt{T}}\text{ }\frac{\epsilon}{v_{z}}\right)}$, where the cell length, $l$, is dependent on the absolute pressure, $p$, according to Eq. 11. Consequently, the transmission can be plotted as a function of ${\left(\frac{lp}{\sqrt{T}}\right)}$. By this choice, one expects a close-to-exponential dependence of the measured transmissions on ${\left(\frac{lp}{\sqrt{T}}\right)}$, which also includes the broadening due to the dwell time spectrum. These transmission profiles are seen in the top plots of Fig. 4 for D2 gas and Fig. 5 for Ne.
@@
Figure 4: Top: UCN transmission data (black dots) in D2 gas as a function of ${\left(\frac{lp}{\sqrt{T}}\right)}$ (at all measured $T$-s), and the best fits of the two alternative models with example values for the independent variable $T$ in case of YK (blue=100 K, yellow=300 K, red=400 K lines), and one line with LETA (black). Center: Point-by-point residuals of the YK and LETA model fits to the measured data, indicating smaller deviations for the YK model fit, showing data at all measured temperatures in a color scale. Bottom: Histograms of the above residuals showing a much narrower and symmetrical distribution for the case of the YK model.
@@
In contrast to D2, the Ne data fits with different temperatures shown in Fig. 5 converge with a slope in $T$ compatible with zero, see also Table 2, hence the fit involving Eq. 6 reproduces the LETA model.
@@
The fit results are summarized in Table 2. As in the results for other gases (Table 1), the uncertainties of $\epsilon_{0}$ are for both gases, D2 and Ne, dominated by the systematic contribution from $N_{\mathrm{dwell}}$, whereas, for $\epsilon_{T}$ the uncertainties are dominated by statistics, as the latter is a relative parameter. The neutron counts per measurement point for both detectors were above $3\times 10^{5}$. However, considering only pure neutron count rate statistics would neglect other effects, such as minor changes in the neutron spectrum due to slow changes in the neutron production efficiency of the source.
@@
Species $\epsilon_{0}$ stat. syst. $\epsilon_{T}$ stat. syst.
(barn $(\frac{\text{barn }m/s}{K^{1/2}})$$\pm$ $\pm$ $10^{-6}/K$$\pm$ $\pm$
Ne 100.6 1.0 11.0 54 38 10
D2 1238.8 3.8 66 559 10 18
@@
Returning to Table 1, the absolute value of the D2 total cross section for UCNs is given under ‘literature’ as calculated by integrating Eqs. (A4) and (A5) in Ref. [16] and adding in a 2:1 ortho-para RT-equilibrium ratio. Our measured value for RT is smaller than reported in [23] by a factor of 1.2 of the quadratically combined error bars. In Fig. 6 we plotted the D2 total cross section as a function of temperature extracted by applying the YK model to the measured data using the parameters in Table 2. We also measured the total cross section of UCNs in H2 gas. The data were analyzed by applying the Young-Koppel model, however, in a more limited temperature range (250-350 K) and thus with less sensitivity to the details of this model. The theoretical YK differential cross sections, Eqs. (29)-(30) in Ref. [16], for para- and ortho-hydrogen were integrated over final energy and solid angle, and added in a 1:3 para to ortho equilibrium ratio at RT. The measurement result for H2 and the theoretical expectation are in very good agreement, see Table 1, consistent within the systematic uncertainties. Our extracted value for RT is smaller than reported in [24] by a factor of 1.1 of the combined error bars.
Figure 6: Temperature dependence of the D2 total cross section applying the YK model to all measured data. To emphasize the difference, the dashed line indicates the $\sqrt{T}$ dependence of the monoatomic approximation.